3) What work is done by a force F=(2i+4j)N, that moves a particle from a position ri=(3i+4j)m to a position rf=(-4i-2j)m?
1
Expert's answer
2019-08-07T10:00:15-0400

The work is given by formula


"W=\vec{F}×\vec{s} (1)"

In our case, we have


"\vec{s} =(-4i-2j) - (3i+4j)=-7i-6j (2)"

We put (2) in (1)


"W=(2i+4j)( -7i-6j) = -14i - 24j=-38"

Answer

-38 J